Fill in the blanks.
(a) (-8) + …. = 0
(b) 13 + …. = 0
(c) 12 + (-12) = ….
(d) (-4) + …. = – 12
(e) …. -15 = – 10.

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
(a) (-8) + (additive inverse of -8) = 0 = (-8) + (8) = 0 ∴ Value of blank is 8 (b) 13 + (additive inverse of 13) = 0 = 13 + (-13) = 0 ∴ Value of blank is – 13 (c) 12 + (-12) = 0 [∵ -12 is additive inverse of 12] ∴ The Value of blank is 0 (d) (-4) + (-8) = -[4 + 8] = -12 ∴ Value of blank is -8. (e) (+5) – 15 = -10 ∴ Value of blank is +5